OpenStudy (anonymous):

A car has a kinetic energy of 4.32×10 5 J when traveling at a speed of 25 m/s. What is its mass? Answer in units of kg

OpenStudy (anonymous):

1/2mv^2= 4.32*10^5 mv^2 = 2.16*10^5 625m= 2.16*10^5 m= 354.6 kg